OpenStudy (anonymous):

A number is divided by 3. If 8 is subtracted from the quotient you get -6.

OpenStudy (anonymous):

number is 2

OpenStudy (anonymous):

A. -6 B. 3 C. 6 D. 42

OpenStudy (anonymous):

but how come 2 is divided by 3??

OpenStudy (anonymous):

let the number be x so what we have is if 8 is subtracted from number then we have the new number x-8 and when it is divided by 3 it gives -6 as quotient ( i assume this is what u said ) therefore \[\frac{ x-8 }{3} = -6\]

OpenStudy (anonymous):

Hey!! @SandeepReddy

OpenStudy (anonymous):

now solve for x!

OpenStudy (anonymous):

Yes LizzyLove<3

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

Its Okay Kanwar

OpenStudy (anonymous):

Its Okay Pa ji ! :P

OpenStudy (phi):

A number is divided by 3 I would call the number x this says x/3 (it gives a quotient) 8 is subtracted from the quotient that would be written x/3- 8 you get -6. that means x/3 -8 = -6 can you finish ?

OpenStudy (anonymous):

Nothing just saying hey!!

OpenStudy (anonymous):

@SandeepReddy